Grade 12Mechanics

The motion of a particle executing S.H.M is given by x=5sin200π(t+1) where x is in metres and time is in seconds. Find the time period of oscillation.

Priya
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
from equation, omega = angular frequency = 200*pi = 2*pi/T

 

=>Time period ,  T = 2/200 = 0.01 seconds

 

Hope it helps